Problem 1) A Fantastic Transportation Company bought today a new Tractor unit for use over the road for shipping products all over the USA for $98,000. And this Tractor unit requires an additional working capital of $3,000 to install a GPS tracking system. It is estimated that the Salvage Value of tractor unit is going to be equal to $8,000 at the end of its useful life. Fantastic Company uses an After-Tax MARR of 15%. Calculate: ONLY the depreciation amounts for:

A) All the depreciation amounts for this Tractor over its useful life by using the 200% Declining-Balance method.

B) All the depreciation amounts for this Tractor over its useful life by using the Straight-Line method.

C) All the depreciation amounts for this Tractor over its useful life by using the MACRS-GDS method

D) All the depreciation amounts for this Tractor over its useful life by using the MACRS-ADS method.

Kindly use the following table to fill the required amounts after you have shown ALL the equations and calculations below it in order to answer the above questions. Fill only the required columns and rows.

EOY 200% Declining-Balance method Depreciation Straight-line method Depreciation MACRS-GDS method Depreciation MACRS-ADS method Depreciation

Please show calculations for the methods, and please no Excel usage. Thank you.
Answer #1

A) Useful life = 10

Straight line depreciation percent = 1/10 = 0.1 or 10% per year

Depreciation rate = 10% * 2 = 20% per year

Depreciation for each year= Rs. 101,000 * 20% = Rs. 20,200

B) Annual Depreciation expense = (98,000+3000-8,000) / 10 = Rs. 9,300